Let h(u,v)=ln⁡(n-u^2+v)+√(√v-u)  , Where n=557. Find its domain and sketch the Graph


1
Expert's answer
2021-05-07T09:38:31-0400

We have given the function,

"h(u,v) = log(n-u^2+v) + \sqrt{{\sqrt{v-u}}}" where "n = 557"

Domain of this function is

{"(u,v) \isin R^2 : u \le \sqrt v \hspace{3mm}and \hspace{3mm} u^2< v+557 \hspace{3mm}and \hspace{3mm}v \ge 0" }

Range of the function is

{ "y \isin 6.88708 \le y \le 9.699" }